### v2

@@ Abstract
We investigate the sensing capacity of non-equilibrium dynamics in quantum systems exhibiting Bloch oscillations. By focusing on the resource efficiency of the probe, quantified by quantum Fisher information, we find different scaling behaviors in two different phases, namely localized and extended. Our results provide a quantitative ansatz for quantum Fisher information in terms of time, probe size, and the number of excitations. In the long-time regime, the quantum Fisher information is a quadratic function of time, touching the Heisenberg limit. The system size scaling drastically depends on the phase changing from super-Heisenberg scaling in the extended phase to size-independent behavior in the localized phase. Furthermore, increasing the number of excitations always enhances the precision of the probe, although, in the interacting systems the enhancement becomes less eminent than the non-interacting probes. This is due to the induced localization by increasing the interaction between the excitations.
@@ 1 Introduction
The delicate nature of quantum systems makes them naturally suitable for sensing gravitational, magnetic, and electric fields with unprecedented precision well beyond the capacity of classical probes [1]. In order to estimate an unknown parameter $h$, encoded in the quantum state of a probe $\rho(h)$, one has to perform measurement. For a given measurement, described by a set of projectors $\{\pi_{x}\}$, the results follow a classical probability distribution in which any outcome appears with the probability $p_{x}(h)=\mathrm{Tr}[\pi_{x}\rho(h)]$. In this situation, the precision of the probe for sensing parameter $h$, quantified by standard deviation $\delta h$, is fundamentally bounded by Cramér-Rao inequality $\delta h{\geq}1/\sqrt{\mathcal{M}\mathcal{F}_{C}}$, where $\mathcal{M}$ is the number of samples and $\mathcal{F}_{C}$ is called Classical Fisher information (CFI) which is defined as [2, 3, 4]
$\mathcal{F}_{C}(h)=\sum_{x}\frac{1}{p_{x}(h)}(\frac{dp_{x}}{dh})^{2}.$ (1)
In order to obtain a measurement-independent quantity, one can maximize the CFI with respect to all possible measurements to obtain Quantum Fisher Information (QFI), namely $\mathcal{F}_{Q}=\underset{\{\pi_{x}\}}{\mathrm{max}}\mathcal{F}_{C}$. The QFI indeed describes the ultimate precision limit for which the Cramér-Rao inequality can be written as
$\delta h\geq\frac{1}{\sqrt{\mathcal{M}\mathcal{F}_{C}}}\geq\frac{1}{\sqrt{\mathcal{M}\mathcal{F}_{Q}}},$ (2)
Interestingly, the maximization over all possible measurements in the definition of the QFI can be reduced to a closed formula [3]. For instance, in the case of pure states $\rho(h){=}\mathinner{|{\Psi(h)}\rangle}\mathinner{\langle{\Psi(h)}|}$ the QFI takes the form [3]
@@
The performance of a sensor is determined by the scaling of its QFI with respect to resources (probe size and time), namely $\mathcal{F}_{Q}{\sim}L^{\beta}$, where $L$ is the resource and $\beta$ is an exponent. While in the absence of quantum features, one can at best achieve $\beta{=}1$ (i.e. standard limit), in quantum probes, the precision might be enhanced to $\beta>1$, known as quantum enhanced sensitivity, in which a special case is $\beta{=}2$ (i.e. Heisenberg limit) [5].
@@
Quantum criticality is known to be instrumental for achieving quantum enhanced sensitivity. In fact, various forms of criticality have been exploited for sensing purposes, including first-order [6, 7, 8], second-order [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25], dissipative [26, 27, 28, 29, 30, 31, 32, 33, 34, 35], topological [36, 37, 38, 39], Floquet [40, 41], time crystals [42, 43, 44], non-Hermitian systems [45, 46, 47, 48, 49, 50], and Stark [51, 52, 53] phase transitions. However, despite a few attempts for realizing criticality-based sensors [54, 20], their experimental implementation faces real challenges. Normally, in most of such sensors, the probe should be initialized in one of its eigenstates, e.g. the ground state, which is very challenging in practice and may require extra resources to be achieved [19]. In addition, the parameter interval over which such criticality-based probe achieves quantum enhanced sensitivity is often very narrow covering only around the phase transition point. A possible solution for addressing these challenges is to go beyond equilibrium physics and exploit non-equilibrium dynamics of many-body systems [55, 56, 57, 58, 59], which is easy to implement in various physical platforms [60, 61, 62, 63, 64]. Such systems do not require complex initialization and may allow quantum enhanced sensitivity over a wider range. Several open questions exist, including: (i) what types of non-equilibrium systems may lead to quantum enhanced sensitivity? and (ii) can such probes operate optimally across an entire phase of matter rather than a narrow region around the phase transition point?
@@
Bloch oscillation is a fundamental phenomenon in condensed matter physics in which a particle under the impact of a gradient potential (i.e. a constant force) oscillates in a regular lattice [65]. The gradient potential naturally makes the neighboring lattice sites off-resonant and hence suppresses the tunneling of particles. Unlike classical systems in which particles simply move in the direction of the external force, this suppression of tunneling in quantum systems induces Wannier-Stark localization in the system [66]. The localization properties of Wannier-Stark systems have been extensively studied in both theory [67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90] and experiments [91, 92, 93, 94]. Bloch oscillations have also been observed in semiconductor devices [95], optical waveguides [96, 97, 98], cold atoms in optical lattices [99] and superconducting simulators [100, 101]. Unlike the conventional second-order quantum phase transitions which only affect the ground state, the Wannier-Stark localization takes place across the entire spectrum. This implies that the impact of such transition should be observable in non-equilibrium dynamics, where many eigenstates are involved. Recently, it has been shown that individual eigenstates of the Wannier-Stark systems can be used for sensing the gradient field with super-Heisenberg precision [51]. A natural question is whether one can exploit experimental-friendly non-equilibrium dynamics of such systems for sensing purposes.
@@
In this paper, we explore the sensing capacity of the Bloch oscillations in single- and many-body Stark systems. We provide a comprehensive analysis for the scaling of the QFI in terms of time, probe size, and the number of excitations. We show that indeed many-body systems with Bloch oscillations may allow for sensing precision with super Heisenberg scaling. Unlike critical probes at equilibrium we do not demand complex initialization and the probe operates optimally over the entire extended phase.
@@ 2 Bloch oscillation in Stark systems
where $J$ is the exchange coupling, $h$ is the gradient field and $\mathinner{|{l}\rangle}$ represents the excitation at site $l$. The gradient field induces a position-dependent energy shift on each site which suppresses the particle tunneling. This leads to the localization of the wave function into a limited number of sites, known as Wannier-Stark localization [66]. In the limit of $L{\rightarrow}\infty$, the eigenstates of Hamiltonian (4), the Wannier-Stark states, are found as $\mathinner{|{E_{m}}\rangle}=\sum_{l=1}^{L}\mathcal{J}_{l-m}\big{(}\frac{2J}{h}\big{)}\mathinner{|{l}\rangle}$ whose corresponding eigenvalues are $E_{m}=mh$ [67]. Here, $\mathcal{J}_{l}(\cdot)$ are the Bessel function of the first kind. The state $\mathinner{|{E_{m}}\rangle}$ is a superposition of the states centered at the $m$th site and spreading to the neighboring sites with the extend about $4J{/}h$ lattice periods [102]. Any initial state $\mathinner{|{\Psi(0)}\rangle}=\sum_{l}f_{l}\mathinner{|{l}\rangle}$, evolves under the action of the Hamiltonian $\mathinner{|{\Psi(t)}\rangle}=e^{-iHt}\mathinner{|{\Psi(0)}\rangle}$ which can be cast into $\mathinner{|{\Psi(t)}\rangle}=\sum_{l,l^{\prime}=1}^{L}\mathcal{K}_{l,l^{\prime}}(t)f_{l^{\prime}}\mathinner{|{l}\rangle}$ with the evolution propagator $\mathcal{K}_{l,l^{\prime}}(t){=}\sum_{m=1}^{L}\braket{l}{E_{m}}\!\braket{E_{m}}{l^{\prime}}e^{-iE_{m}t}$ [67]. Note that in the large $L$ limit, the energies are equidistant, i.e. $\Delta E{\equiv}E_{m+1}{-}E_{m}{=}h$. Hence, the evolution propagator exhibits a time-periodic behavior with the characteristic time $T_{\text{Bloch}}=2\pi/h$, which leads to the Bloch oscillations. In order to see the dynamical behavior of the system, we depict $P_{l}(t){=}|\langle{l}|\Psi(t)\rangle|^{2}$ as a function of time in Figs. 1(a)-(c) for three different values of $h$ in a system of size $L{=}100$, initialized at the central site. For larger values of $h$, shown in Fig. 1(a), one clearly observes the periodic spreading of the excitation over a finite distance and returning to its original site. Here, the Bloch oscillations indicate a localized phase. By decreasing $h$ the extent of periodic wave packet spreading increases until $h{=}h_{c}{\simeq}8J{/}L$, at which the wave packet delocalizes across the entire system and then localizes back, see Fig. 1(b). This is the point of transition in which the localized phase transforms into an extended phase. Indeed, by further decreasing of $h$, a new behavior emerges as the excitation spreads across the entire system but does not fully return to its initial localized site, signaling an ergodic behavior [102, 51].
@@ 3 Model 1: Single excitation probe
Figure 2: (a) Long-time behaviour of $\mathcal{F}_{Q}/t^{2}$ as a function of $h/J$ for different system sizes. (b) The scaling of the long-time limit of $\mathcal{F}_{Q}/t^{2}$ with respect to the system size $L$ at the transition point $h{=}h_{c}$ (green squares) and deep into the extended phase $h/J{=}0.01$ (blue triangles). The dashed lines show the best fitting functions. (c) Exponent $\beta$ as a function of $Lh/J$ for various system sizes $L$. The transition point is specified at $h_{c}=8J/L$ whose corresponding exponent is $\beta{=}2$.
After investigating the time dependence of the QFI, we now focus on scaling with respect to the probe size $L$. In Fig. 2(a) we plot the long-time values of $\mathcal{F}_{Q}/t^{2}$ as a function of $h$ for different system sizes. The localization transition points for different system sizes $L$ are clearly identified, which happens around $h_{c}{\simeq}8J/L$. In the localization phase, $\mathcal{F}_{Q}$ becomes size-independence and scales inversely with square of the gradient field. On the other hand, the QFI becomes size-dependent for $h\leq h_{c}$. These suggest that
@@
where the scaling exponent $\beta$ may vary with $h$. Indeed, at the transition point $h_{c}{\simeq}8J/L$, the QFI from both sides of the transition point, given in Eqs. (5), match with each other which results in $\beta{=}2$. This is confirmed numerically as for different system sizes we find maximized $\mathcal{F}_{Q}/t^{2}$ with respect to $h$ at large time $t$ where $\mathcal{F}_{Q}/t^{2}$ is saturated, see the green squares in Fig. 2(b). Interestingly, in the extended phase, the scaling becomes even better as for $h/J{=}0.01$ the exponent $\beta$ increases to $\beta\simeq 2.5$, depicted by blue triangles in Fig. 2(b). Finally, to better understand the metrological aspects of our model we extract the scaling exponent $\beta$ for arbitrary values of $h$, see Appendix A for more details. The results are shown in Fig. 2(c), where one can see a clear super-Heisenberg scaling with $\beta>2$ in the entire extended phase. The maximum scaling is obtained around $hL/J{\simeq}5$ whose exponent is $\beta{\simeq}3.5$. As mentioned before, at the transition point $h{=}h_{c}$ the QFI scales with Heisenberg precision, i.e. $\beta{=}2$, which then drops to $\beta{=}0$ in the localized phase.
(a)Refer to caption(b)
Refer to caption(c)Refer to caption(d)

### v3

@@ Abstract
We investigate the sensing capacity of non-equilibrium dynamics in quantum systems exhibiting Bloch oscillations. By focusing on the resource efficiency of the probe, quantified by quantum Fisher information, we find different scaling behaviors in two different phases, namely localized and extended. Our results provide a quantitative ansatz for quantum Fisher information in terms of time, probe size, and the number of excitations. In the long-time regime, the quantum Fisher information is a quadratic function of time, touching the Heisenberg limit. The system size scaling drastically depends on the phase changing from quantum-enhanced scaling in the extended phase to size-independent behavior in the localized phase. Furthermore, increasing the number of excitations always enhances the precision of the probe, although, in the interacting systems the enhancement becomes less eminent than the non-interacting probes. This is due to the induced localization by increasing the interaction between the excitations. We show that a simple particle configuration measurement together with a maximum likelihood estimation can closely reach the ultimate precision limit in both single- and multi-particle probes.
@@ 1 Introduction
The delicate nature of quantum systems makes them naturally suitable for sensing gravitational, magnetic, and electric fields with unprecedented precision well beyond the capacity of classical probes [1]. In order to estimate an unknown parameter $h$, encoded in the quantum state of a probe $\rho(h)$, one has to perform measurement. For a given measurement, described by Positive Operator-Valued Measure operators $\{\pi_{x}\}$ (POVM), the results follow a classical probability distribution in which any outcome appears with the probability $p_{x}(h){=}\mathrm{Tr}[\pi_{x}\rho(h)]$. By post-processing the measurement results, one can construct an estimator $\hat{h}$. The estimator has to be asymptotically unbiased in the sense that its expectation value gives the real value of the unknown parameter, namely $\langle\hat{h}\rangle{=}h$, in the limit of large measurement samples. In this situation, the precision of the probe for sensing parameter $h$ is quantified by the standard deviation of the estimator, which we call it $\delta h$. For a given measurement $\{\pi_{x}\}$ and an unbiased estimators, the precision $\delta h$ is asymptotically bounded by Cramér-Rao inequality $\delta h{\geq}1/\sqrt{\mathcal{M}\mathcal{F}_{C}}$, where $\mathcal{M}$ is the number of samples and $\mathcal{F}_{C}$ is called Classical Fisher information (CFI) which is defined as [2, 3, 4, 5]
$\mathcal{F}_{C}(h)=\sum_{x}\frac{1}{p_{x}(h)}\left(\frac{dp_{x}}{dh}\right)^{2}.$ (1)
The estimator that can saturate the above inequity in the asymptotic limit is called “efficient” estimator. Therefore, in order to saturate the Cramér-Rao inequality of Eq. (1) one has to rely on an optimal estimator which is both unbiased and efficient. Bayesian and maximum likelihood are known for being asymptotically optimal estimators. According to Laplace-Bernstein–von Mises theorem [6, 7], in the limit of large repeating measurement samples, the Bayesian and maximum likelihood converge to the same value of estimation. In particular, estimation by a Bayesian estimator with uniform prior in large repetitions is literally identical to maximum likelihood.
Furthermore, one can go one step further and try to find a precision bound which is independent of measurement basis. To accomplish this, one can maximize the CFI with respect to all possible POVMs to obtain Quantum Fisher Information (QFI), namely $\mathcal{F}_{Q}{=}\underset{\{\pi_{x}\}}{\mathrm{max}}\mathcal{F}_{C}$. The QFI indeed describes the ultimate precision limit for which the Cramér-Rao inequality can be written as
$\delta h\geq\frac{1}{\sqrt{\mathcal{M}\mathcal{F}_{C}}}\geq\frac{1}{\sqrt{\mathcal{M}\mathcal{F}_{Q}}}.$ (2)
While the saturation of the first inequality requires an optimal estimator, the saturation of the second inequality demands both optimized measurement and estimator. Interestingly, the maximization over all possible measurements in the definition of the QFI can be reduced to a closed formula [3], see Refs. [5, 4] for detailed discussions on QFI. For instance, in the case of pure states $\rho(h){=}\mathinner{|{\Psi(h)}\rangle}\mathinner{\langle{\Psi(h)}|}$ the QFI takes the following form [3]
@@
The performance of a sensor is determined by the scaling of its QFI with respect to resources (probe size and time), namely $\mathcal{F}_{Q}{\sim}L^{\beta}$, where $L$ is the resource and $\beta$ is an exponent. In the absence of quantum features, one can at best achieve $\beta{=}1$ (i.e. standard limit). On the other hand, by properly harnessing quantum features, such as entanglement, the precision might be enhanced to $\beta>1$, known as quantum-enhanced sensitivity, in which a special case is $\beta{=}2$ (i.e. Heisenberg limit) [8].
@@
Quantum criticality is known to be instrumental for achieving quantum-enhanced sensitivity. In fact, various forms of criticality have been exploited for sensing purposes, including first-order [9, 10, 11], second-order [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], dissipative [29, 30, 31, 32, 33, 34, 35, 36, 37, 38], topological [39, 40, 41, 42], Floquet [43, 44], time crystals [45, 46, 47], non-Hermitian systems [48, 49, 50, 51, 52, 53], and Stark [54, 55, 56] phase transitions. However, despite a few attempts for realizing criticality-based sensors [57, 23], their experimental implementation faces real challenges. Normally, in most of such sensors, the probe should be initialized in one of its eigenstates, e.g. the ground state, which is very challenging in practice and may require extreme cooling or adiabatic state preparation whose time resource scales undesirably with the system size [22]. In addition, the parameter interval over which such criticality-based probe achieves quantum-enhanced sensitivity is often very narrow covering only around the phase transition point. A possible solution for addressing these challenges is to go beyond equilibrium physics and exploit non-equilibrium dynamics of many-body systems [58, 59, 60, 61, 62], which are easy to implement in various physical platforms [63, 64, 65, 66, 67]. Such systems do not require complex initialization and may allow quantum-enhanced sensitivity over a wider range. Several open questions exist, including: (i) what types of non-equilibrium systems may lead to quantum-enhanced sensitivity? and (ii) can such probes operate optimally across an entire phase of matter rather than a narrow region around the phase transition point?
@@
Bloch oscillation is a fundamental phenomenon in condensed matter physics in which a particle under the impact of a gradient potential (i.e. a constant force) oscillates in a regular lattice [68]. The gradient potential naturally makes the neighboring lattice sites off-resonant and hence suppresses the tunneling of particles. Unlike classical systems in which particles simply move in the direction of the external force, this suppression of tunneling in quantum systems induces Wannier-Stark localization in the system [69]. The localization properties of Wannier-Stark systems have been extensively studied in both theory [70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93] and experiments [94, 95, 96, 97]. Bloch oscillations have also been observed in semiconductor devices [98], optical waveguides [99, 100, 101], cold atoms in optical lattices [102] and superconducting simulators [103, 104]. Unlike the conventional second-order quantum phase transitions which only affect the ground state, the Wannier-Stark localization takes place across the entire spectrum. This implies that the impact of such transitions should be observable in non-equilibrium dynamics, where many eigenstates are involved. Recently, it has been shown that individual eigenstates of the Wannier-Stark systems can be used for sensing the gradient field with quantum-enhanced precision [54]. A natural question is whether one can exploit experimental-friendly non-equilibrium dynamics of such systems for sensing purposes.
@@
In this paper, we explore the sensing capacity of the Bloch oscillations in single- and many-body Stark systems. We provide a comprehensive analysis for the scaling of the QFI in terms of time, probe size, and the number of excitations. We show that indeed many-body systems with Bloch oscillations may allow for sensing precision with quantum-enhanced scaling. Unlike critical probes at equilibrium we do not demand complex initialization and the probe operates optimally over the entire extended phase. We also demonstrate that in a practical setup a simple particle configuration measurement together with a maximum likelihood estimation can closely saturate the bound given by the Cramér-Rao inequality.
@@ 2 Bloch oscillation in Stark systems
where $J$ is the exchange coupling, $h$ is the gradient field and $\mathinner{|{l}\rangle}$ represents the excitation at site $l$. The gradient field induces a position-dependent energy shift on each site which suppresses the particle tunneling. This leads to the localization of the wave function into a limited number of sites, known as Wannier-Stark localization [69]. In the limit of $L{\rightarrow}\infty$, the eigenstates of Hamiltonian (4), the Wannier-Stark states, are found as $\mathinner{|{E_{m}}\rangle}{=}\sum_{l=1}^{L}\mathcal{J}_{l-m}\big{(}\frac{2J}{h}\big{)}\mathinner{|{l}\rangle}$ whose corresponding eigenvalues are $E_{m}{=}mh$ [70]. Here, $\mathcal{J}_{l}(\cdot)$ are the Bessel function of the first kind. The state $\mathinner{|{E_{m}}\rangle}$ is a superposition of the states centered at the $m$th site with the spreading width about $4J{/}h$ lattice periods [105]. Any initial state $\mathinner{|{\Psi(0)}\rangle}{=}\sum_{l}f_{l}\mathinner{|{l}\rangle}$, evolves under the action of the Hamiltonian $\mathinner{|{\Psi(t)}\rangle}{=}e^{-iHt}\mathinner{|{\Psi(0)}\rangle}$ which can be cast into $\mathinner{|{\Psi(t)}\rangle}{=}\sum_{l,l^{\prime}=1}^{L}\mathcal{K}_{l,l^{\prime}}(t)f_{l^{\prime}}\mathinner{|{l}\rangle}$ with the evolution propagator $\mathcal{K}_{l,l^{\prime}}(t){=}\sum_{m=1}^{L}\braket{l}{E_{m}}\!\braket{E_{m}}{l^{\prime}}e^{-iE_{m}t}$ [70]. Note that in the large $L$ limit, the energies are equidistant, i.e. $\Delta E{\equiv}E_{m+1}{-}E_{m}{=}h$. Hence, the evolution propagator exhibits a time-periodic behavior with the characteristic time $T_{\text{Bloch}}{=}2\pi/h$, which leads to the Bloch oscillations. In order to see the dynamical behavior of the system, we depict $P_{l}(t){=}|\langle{l}|\Psi(t)\rangle|^{2}$ as a function of time in Figs. 1(a)-(c) for three different values of $h$ in a system of size $L{=}100$, initialized at the central site. For larger values of $h$, shown in Fig. 1(a), one clearly observes the periodic spreading of the excitation over a finite distance and returning to its original site. Here, the Bloch oscillations indicate a localized phase. By decreasing $h$ the extent of periodic wave packet spreading increases until $h{=}h_{c}{\simeq}8J{/}L$, at which the wave packet delocalizes across the entire system and then localizes back, see Fig. 1(b). This is the point of transition in which the localized phase transforms into an extended phase. Indeed, by further decreasing of $h$, a new behavior emerges as the excitation spreads across the entire system but does not fully return to its initial localized site, signaling an ergodic behavior [105, 54].
@@ 3 Model 1: Single excitation probe
Figure 2: (a) Long-time behavior of $\mathcal{F}_{Q}/t^{2}$ as a function of $h/J$ for different system sizes. (b) The scaling of the long-time limit of $\mathcal{F}_{Q}/t^{2}$ with respect to the system size $L$ at the transition point $h{=}h_{c}$ (green squares) and deep into the extended phase $h/J{=}0.01$ (blue triangles). The dashed lines show the best fitting functions. (c) Exponent $\beta$ as a function of $Lh/J$ for various system sizes $L$. The transition point is specified at $h_{c}{=}8J/L$ whose corresponding exponent is $\beta{=}2$.
After investigating the time dependence of the QFI, we now focus on scaling with respect to the probe size $L$. In Fig. 2(a) we plot the long-time values of $\mathcal{F}_{Q}/t^{2}$ as a function of $h$ for different system sizes. The localization transition points for different system sizes $L$ are clearly identified, which happens around $h_{c}{\simeq}8J/L$. Note that this is the point at which the Bloch oscillations cover the entire system. In the localization phase, $\mathcal{F}_{Q}$ becomes size-independent and scales inversely with square of the gradient field. On the other hand, the QFI becomes size-dependent for $h\leq h_{c}$. These suggest that
@@
where the scaling exponent $\beta$ may vary with $h$. Indeed, at the transition point $h_{c}{\simeq}8J/L$, the QFI from both sides of the transition point, given in Eqs. (5), match with each other which results in $\beta{=}2$. This is confirmed numerically as for different system sizes we find maximized $\mathcal{F}_{Q}/t^{2}$ with respect to $h$ at large time $t$ where $\mathcal{F}_{Q}/t^{2}$ is saturated, see the green squares in Fig. 2(b). In the extended phase, the scaling becomes even better as for $h/J{=}0.01$ the exponent $\beta$ increases to $\beta\simeq 2.5$, depicted by blue triangles in Fig. 2(b). Finally, to better understand the metrological aspects of our model we extract the scaling exponent $\beta$ for arbitrary values of $h$, see Appendix A for more details. The results are shown in Fig. 2(c), where one can see a clear quantum-enhanced scaling with $\beta>2$ in the entire extended phase. The maximum scaling is obtained around $hL/J{\simeq}5$ whose exponent is $\beta{\simeq}3.5$. As mentioned before, at the transition point $h{=}h_{c}$ the QFI scales with Heisenberg precision, i.e. $\beta{=}2$, which then drops to $\beta{=}0$ in the localized phase. It is worth emphasizing that the region below the critical point, where $\beta{>}2$ is a region which only exit is finite system sizes and its width shrinks by increasing the length of the chain. Therefore, at the large $L$ limit, where $h_{c}{\rightarrow}0$, the obtainable exponent would be $\beta{\leq}2$.
The above discussion and the studies in the following sections rely on the scaling of $\mathcal{F}_{Q}/t^{2}$ at long times. However, we have to specify what long time means. As mentioned above, in the localized phase the Bloch oscillations occur with the period of $T_{\mathrm{Bloch}}=2\pi/h$. By careful checking of the evolutions in Fig. 1(d), one can see that $\mathcal{F}_{Q}/t^{2}$ saturates only after a few ($\sim 5$) oscillations. By decreasing $h$ the system becomes less localized and more extended which makes the saturation time longer. Nonetheless, the saturation time in the units of $T_{\mathrm{Bloch}}$ decreases. The behavior remains the same in the extended phase such that the saturation takes place after $t{=}2\pi/h$. Therefore, generally speaking, the long-time behavior can be obtained after a few ($\sim 1-5$) multiples of $t{=}2\pi/h$ for any values of the gradient field $h$.